OpenStudy (anonymous):

How do you calculate the explicit solution to dP/dt=kP-h? (k and h are positive constants) How do you calculate the explicit solution to dP/dt=kP-h? (k and h are positive constants) @Mathematics

OpenStudy (amistre64):

this is a first order linear diffy q right?

OpenStudy (amistre64):

P' - kP = -h is the usual set up, but im cant recall to well what an explicit solution refers to

OpenStudy (amistre64):

multiply both sides by e^{-kt} P'e^(-kt) - k e^(-kt)P = -h e^(-kt) this creates a situation where the left side is the result of a product rule from derivatives int (P'e^(-kt) - k e^(-kt)P = -h e^(-kt)) dt e^(-kt)P = -h * int (-h e^(-kt)) dt

OpenStudy (amistre64):

this familiar? or do we do separation of variables ?

OpenStudy (amistre64):

e^(-kt)P = -h * int (e^(-kt)) dt e^(-kt)P = -h * (e^(-kt)/-k + C) e^(-kt)P = -he^(-kt)/-k - hC P = h/k - hC e^(kt) is this result a general or explicit? i cant recall

OpenStudy (amistre64):

\[dP=(kP-h) dt\] \[\frac{1}{(kP-h)}dP= dt\] \[\int\frac{1}{(kP-h)}dP= \int dt\] \[\frac{1}{k}\int\frac{k}{(kP-h)}dP= t + C\] \[\frac{1}{k}ln|kP-h|= t + C\] \[ln|kP-h|= kt + kC\] \[exp(ln|kP-h|= kt + kC\] \[kP-h= e^{kt} + C\] \[kP= e^{kt} + C+h\] \[P= \frac{e^{kt}}{k} + C\] maybe

OpenStudy (amistre64):

http://www.wolframalpha.com/input/?i=dP%2Fdt%3DkP-h the wolf likes my first getup

OpenStudy (anonymous):

That looks right. Many thanks!

OpenStudy (amistre64):

yep, i get to take diffy qs next term :)

OpenStudy (anonymous):

You must be some kind of genius if you can do that and not have diffy qs!

OpenStudy (amistre64):

genius maybe :) but i see an algebra mistake on this last bit \[ln|kP-h|= kt + kC\] \[e(ln|kP-h|= kt + kC)\] \[kP-h= e(kt + kC)\] \[kP-h= e^{kt} * e^{kC}\] \[kP-h= c_1e^{kt}\] \[kP= c_1e^{kt}+h\] \[P= c_1e^{kt}+\frac{h}{k}\] so it comes out them same; the constant on the "e" tends to absorb any constants that get near it